# CGRA4ML: A Framework to Implement Modern Neural Networks for Scientific Edge Computing

### v2

@@ VI Results
[…]
Figure 7: Scaling of hardware resource and power consumption on a Xilinx ZCU104 FPGA, for different bitwidths and PE columns, with PE rows = 8
Figure 8: Performance analysis of ResNet-50 on different CGRA array sizes. Performance efficiency measures how well each layer is dynamically mapped over statically configured hardware. While larger arrays reduce off-chip data movement due to high data reuse, they have lower utilization for smaller layers. The unified dataflow of the array (Sec. IV-B) maximizes the utilization to near 100% and minimizes data movement for most layers.
Model ResNet-50 PointNet
Bits 4 4
PEs (7,96) (32,32)
Frequency (MHz) 250 250
FFs 101706 69277
LUTs 82200 100076
BRAMs 6 4.5
Static Power (W) 0.700 0.700
Dynamic Power (W) 3.847 3.840
Total Power (W) 4.547 4.540
GOPs/W 37.3 56.8
TABLE III: Implementation of ResNet-50 and Pointnet on ZCU104 FPGA
VI-A ResNet-50
@@
ResNet-50 [49] is a deep convolutional neural network with 5.36 billion operations in 53 convolutional layers, one dense layer, two pooling layers, and 17 skip connections. Its input is a $224{\times}224{\times}3$ tensor, and outputs are 1000 classes, and it is run on an array of $7{\times}96{=}672$ PEs.
@@
TABLE IV: Comparing Jet-tagger model on CGRA4ML with different array sizes and on hls4ml with varying reuse factors
VI-B PointNet & Jet Tagger
PointNet [50] is a model developed for particle classification as a stack of three 1D convolutions, three dense layers, and two pooling layers. An input size of 10.6K requires 1.18 billion operations for a batch size of 16. We demonstrate this on an array of size $16{\times}32{=}512$. Jet Tagger is a small model with 117K operations in 4 dense layers, widely used for benchmarking. Table IV compares the resource and power consumption of hls4ml and CGRA4ML for the Jet tagger model.
VI-C ASIC Implementation
Technology 7nm 65nm
PEs 768 192 192
Frequency (MHz) 1000 1000 250
Area (mm2) 0.164 0.049 0.47
Static power (mW) 23.62 7.38 54.04
Dynamic power (mW) 381.1 119.9 26.80
Total power (mW) 404.7 127.2 80.84
GOPs/W 1898 1509 594
TABLE V: ASIC results from Cadence toolflow with 7nm and 65nm TSMC LP CMOS PDKs for CGRA configurations with $8{\times}24$ and $8{\times}96$ rows and columns generated using CGRA4ML
@@ VI-C ASIC Implementation
Figure 9: CGRA of $R{\times}C{=}8{\times}24$ configuration generated by CGRA4ML, synthesized with Cadence Genus and placed & routed with Cadence Innovus using TSMC 65nm LP CMOS PDK, being taped out with NanoSoC platform.
As described in Sec. V-E2, we are collaborating with Arm Research to tape out an SoC built through CGRA4ML. We utilize NanoSoC[48], an open-source, ARM-based SoC generation platform built in collaboration with ARM Research. Our CGRA is integrated with the ARM processing system through the DMA360 open-source IP developed and provided in the NanoSoC framework that connects three AXI-Stream interfaces to the AHB interfaces of the NanoSoC interconnect. The ASIC generation toolflow of CGRA4ML is demonstrated through the results presented in Table V. The CGRA from CGRA4ML delivers high performance in a few milliwatts, making it suitable for scientific computing applications at the edge. Our $8{\times}24$ variant of the CGRA with 192 PEs, generated by CGRA4ML, is being taped out into an SoC with the TSMC 65nm LP CMOS PDK.
VII Conclusion
In this paper, we introduced CGRA4ML, a framework for implementing large-scale neural networks in scientific edge computing applications. CGRA4ML successfully bridges the limitations of existing solutions by facilitating off-chip data storage and supporting a diverse array of neural network architectures, including but not limited to Autoencoders, ResNets, and Transformers. Our framework intends to complement hls4ml, a tool widely used by the scientific computing community by handling larger and more complex models that are crucial for advanced scientific computing tasks.
The ability of CGRA4ML to generate RTL designs and runtime firmware, and support end-to-end, vendor-agnostic ASIC and FPGA design flows is a notable advancement, providing a robust solution for researchers and developers aiming to leverage the speed and efficiency of hardware acceleration without the traditional complexity. We demonstrate the effectiveness of our framework by comparing it with hls4ml.
Our contributions are not only technical but also extend to the ease of use and extensibility, thanks to CGRA4ML’s modular and minimal infrastructure. Our Python API, along with its comprehensive verification system, ensures that scientific researchers can focus more on innovation and less on the intricacies of hardware design.
By detailing the framework’s architecture, implementation, and potential for future enhancements, this paper provides a valuable resource for the ongoing development of edge computing technologies. We believe CGRA4ML will play a crucial role in advancing the capabilities of scientific edge computing, paving the way for more sophisticated and real-time data processing at the edge.

### v3

@@ 6.1. Performance Analysis and Practical Guidelines
Operation Percentage
Writeback pointer calculations 47.37%
Writeback 24.34%
Cache flush write 2.14%
Add IT passes 1.10%
Add Bias 4.50%
Activation 6.04%
Residual add 7.50%
Pooling 4.94%
Handshake 1.18%
Table 4. Breakdown of performance penalty of the CPU workload for ResNet-50
6.2. Performance Evaluation of CPU operations
We partition the workload between the CPU and the CGRA as shown in Fig. 5. Lightweight, pointwise operations are done in the CPU, while heavy, parallelizable operations are done through our CGRA. While this partition allows us implement additional pointwise operations and activations as required, with minimal modification of hardware, our evaluations show that often CPU operations become the bottleneck in addition to data transfer between the CPU and the custom hardware. Depending on the workload, over 50% of the execution time is spent on CPU operations, and Table 4 demonstrates the breakdown of execution time in the CPU for ResNet-50. As observed, we can achieve significant speedup by partially parallelizing the most compute-intensive operations, such as calculating writeback pointers and packing writeback words on custom hardware. We are currently working on this to achieve the performance boost without sacrificing extensibility of the bundle architecture. Our fusion strategy includes writing an extensive set of bundle parameters to a register bank at startup, breaking the CGRA-CPU dual execution into a three-stage flow of CGRA-CPU-Writeback with locks and handshaking for synchronization. The new writeback stage would increase the performance by packing output bits, calculating all pointers in parallel and performing writes without CPU intervention.
6.3. Micro-benchmarking
Refer to caption Synthetic workloads of cascading dense layers
Figure 13. Synthetic workloads of dense autoencoder models with 4-32 layers (x-axis), where each dense layer has 16 inputs and 16 outputs and ReLU activation is used between the layers. The model is quantized to have 8-bit inputs, weights, and biases, which are randomized with a normal distribution. Inference is performed with a batch size of 32. The CGRA is configured to have 16 rows, 32 columns, an accumulator width of 24 bits, and 128-bit wide AXI ports. hls4ml stores weights on-chip due to its dataflow architecture, while cgra4ml stores weights off-chip, but buffers them in ping-pong buffers to reuse across the inference of a layer. We do the benchmarking on the ZCU104 development board. cgra4ml maintains flat resource usage, while hls4ml-generated hardware shows linear growth with the number of layers. We report steady-state inverse-throughput, which is the II (initiation interval) for hls4ml. Latency increases linearly for both, with hls4ml showing a slower rate.
We first conduct a micro-benchmarking experiment to observe how the resource usage and latency of hardware generated by hls4ml and cgra4ml scale with neural network size. The synthetic workload is a stack of dense layers with ReLU between layers; each layer is $16{\times}16$ with bias. All tensors (inputs, weights, biases) are 8-bit quantized; weights and biases are initialized from a zero-mean normal distribution. We report the time to process one batch of 32 inputs. The cgra4ml CGRA uses a $16{\times}32$ PE array with 24-bit accumulators and communicates via 128-bit AXI. On ZCU104, hls4ml stores per-layer weights on-chip (consistent with its layer-specialized dataflow), whereas cgra4ml streams weights from off-chip DDR but employs double (ping–pong) buffering to maximize reuse within each layer. Inputs/outputs are transferred over 128-bit DMA, and layer outputs are requantized to 8-bit with round-to-nearest and saturation. The results are summarized in Fig. 13.
The observations align with our discussion in Section 2. hls4ml, a representative example of layer-by-layer hardware generation, produces designs whose resource usage grows linearly with the number of layers. The hardware dataflow is well pipelined, leading to relatively modest growth in latency. In contrast, cgra4ml is a CGRA-based accelerator that reuses the same core across layers, resulting in flat resource consumption. The latency still grows linearly, but limited pipelining and additional dataflow-control overhead lead to a steeper slope. Based on these trends, we identify three regimes by network depth: for small networks (4–8 layers), cgra4ml achieves lower latency (at higher resource usage) due to its high-throughput systolic array; for mid-sized networks (12–24 layers), hls4ml provides both lower latency and resource usage; beyond 28 layers, cgra4ml becomes attractive because of its constant resource footprint, provided the application can tolerate the higher latency relative to layer-specialized designs.
Figure 14. Iso-Resource Microbenchmarking on Pynq-Z2 at clock frequency of 150 MHz and batch size of 16. For cgra4ml, we find a 16×16 PE array (8×8 MAC, 24-bit accumulator) fills the FPGA. For hls4ml, we find the largest dense layer (8×8 MAC) that can fit within the FPGA with no resources exceeding 100% is that with 32 inputs and 64 outputs (2048 multiply-accumulate operations) at reuse factor of 1. We then increase the workload size linearly by 1024 MACs, and find the reuse factor needed to fit that workload in the FPGA using hls4ml. We find that the reuse factor needs to be increased almost exponentially for linear increase in workload size, causing latency and throughput to degrade for large workload sizes.
To complement Fig. 13, we perform an iso-resource benchmarking on a Pynq-Z2 (150 MHz) with 8-bit activations/weights. For hls4ml, we start with a dense layer of 32 inputs and 64 outputs, which we find is the largest dense layer that fits on the FPGA at reuse factor of 1 without any resources exceeding 100%. The workload size is $32{\times}64{=}2048$ multiply-accumulate operations. We then increase the workload size linearly by 1024, finding the reuse factor necessary to fit it into the FPGA. For CGRA4ML, we select the largest array that fits, which is 16×16 PEs with 8-bit × 8-bit multiply and 24-bit accumulation. In our experiment, we find that LUTs are the bottleneck resource. We run them at the same frequency of 150 MHz and batch size of 16. This way, we keep the resource utilization almost constant (100%) and same for cgra4ml and hls4ml throughout the experiment. We report steady-state batch latency and throughput for both designs in Fig. 14 for every workload size.
Increasing hls4ml reuse is often effective for fitting larger layers. But we find that to fit linearly larger workload size (number of multiply-accumulates), the reuse factor needs to be increased exponentially as shown in Fig. 14. This causes latency and throughput to degrade similarly. At high reuse such as 512, hls4ml latency exceeds that of CGRA4ML, which maintains stable latency and higher throughput at the matched resource point. We also found for input size=128 and output size=64, we could not get resource utilization close to 100% even with reuse factor of 8192, which is the largest possible reuse factor for a dense layer with 8192 MAC operations. Therefore, we suggest the users to prefer hls4ml when the model fits with low reuse factor on the target device, and use CGRA4ML when extreme reuse is required to meet resource limits.
6.4. ResNet-50
@@
ResNet-50 (He et al., 2016) is a deep convolutional neural network with 5.36 billion operations in 53 convolutional layers, one dense layer, two pooling layers, and 17 skip connections. Its input is a $224{\times}224{\times}3$ tensor, and outputs are 1000 classes, and it is run on an array of $7{\times}96{=}672$ PEs.
@@
Table 5. Comparing Jet-tagger model on cgra4ml with different array sizes and on hls4ml with varying reuse factors
6.5. PointNet & Jet Tagger
Jet Tagger is a small model with 117K operations in 4 dense layers, widely used for benchmarking. Table 5 compares the Jet Tagger model implemented using cgra4ml with two different CGRA configurations (8×12 and 16×32 PEs) against hls4ml across reuse factors of 1, 16, and 64. Notably, hls4ml achieves lower overall resource usage (FFs and LUTs) and power consumption for this small model. This is expected, as Jet Tagger’s modest compute demand and regular layer structure are well-suited to hls4ml ’s layer-by-layer, statically scheduled architecture, which minimizes hardware overhead when the model fits entirely within the available logic. In contrast, cgra4ml incurs additional overhead from its reconfigurable CGRA infrastructure, runtime firmware, and AXI interfacing, which are optimized for larger, more complex models. For tiny workloads like Jet Tagger, this overhead outweighs the performance benefits of dynamic reconfiguration. Thus, while cgra4ml provides a scalable solution for deeper or irregular networks, hls4ml remains more resource- and power-efficient for simple, compact models, highlighting a natural tradeoff between generality and specialization.
PointNet is a model designed for particle classification. It includes three 1D convolutional layers, three dense layers, and two pooling layers. This results in approximately 1.18 billion operations for a batch size of 16. The model does not synthesize with hls4ml due to its depth, irregular structure, and extensive activation reuse. hls4ml ’s layer-by-layer pipeline architecture requires each layer to be synthesized as a standalone datapath. The entire model must fit into the FPGA fabric simultaneously. Even with reuse factor tuning, hls4ml fails to compile PointNet on the ZCU104 due to resource exhaustion during synthesis and placement. In contrast, cgra4ml supports PointNet by reusing its CGRA fabric across all layers. This enables not only functional deployment but also high energy efficiency (56.8 GOPs/W), making cgra4ml uniquely capable of supporting such mid-sized scientific neural networks that exceed the scalability bounds of traditional HLS-based approaches.
6.6. ASIC Implementation
Technology 7nm 28nm 65nm
Frequency (MHz) 500 500 250
Area (mm2) 0.037 0.175 0.502
Static power (mW) 0.614 25.58 0.130
Dynamic power (mW) 29.8 46.87 80.71
Total power (mW) 30.4 72.45 80.84
GOPs/W 3158 1325 594
Table 6. Power/Performance/Area (PPA) analysis across technology nodes for a CGRA configuration of $R{\times}C=8{\times}24=192$. 28 & 65 nm results use a Cadence Genus/Innovus flow; 7 nm results use a Synopsys DC/ICC2 flow
(a) TSMC 65nm LP (low power)
(b) TSMC 28nm HPC+ (high performance)
(c) TSMC 7nm
(a) 65nm view. (b) 28nm view. (c) 7nm view.
@@ 6.6. ASIC Implementation
Figure 15. CGRA of $R{\times}C{=}8{\times}24$ configuration generated by cgra4ml and taken through the preliminary ASIC scripts. (a) and (b) are synthesized with Cadence Genus and placed & routed with Cadence Innovus, while (c) was synthesized with Synopsys Design Compiler and placed & routed using Synopsys ICC2.
As described in Sec. 5.4, we built a complete system-on-chip (SoC) around the cgra4ml CGRA using NanoSoC (David Flynn, 2023), an open-source ARM-based SoC generation platform developed with ARM Research. The CGRA connects to the ARM Cortex-M subsystem through DMA360, which bridges the accelerator’s three AXI-Stream ports to the AHB interconnect for programmable, high-throughput transfers. Macros for on-chip memories (weights/pixels) are generated with ARM Artisan Memory Compiler, and the SoC is synthesized in Cadence Genus and placed & routed in Cadence Innovus targeting TSMC 28 nm HPCPLUS. This validates that cgra4ml emits silicon-ready RTL that integrates cleanly into a production-style SoC environment.
Beyond first-order synthesis, we executed place-and-route flows across multiple nodes:
  1. (1)
    TSMC 65 nm LP with Cadence Genus/Innovus
  2. (2)
    TSMC 28 nm HPCPLUS with Cadence Genus/Innovus, and
  3. (3)
    TSMC 7 nm with Synopsys DC/ICC2
For 65 nm, we targeted 250 MHz with a custom floorplan that partitions the weight memory into narrow (8-bit) SRAM macros placed symmetrically near AXI-Stream port for weights, and a 12-bit macro for pixels; the resulting design uses 9 SRAM macros. At 28 nm and 7 nm, the SRAM instances are sufficiently compact to avoid width-splitting; these designs employ 3 macros total. Across nodes we used scripted flows from synthesis through P&R, including memory macro integration, floorplanning, clocking, routing, and iterative DRC cleanup.
In addition to the 65 nm and 28 nm implementations above, we evaluated larger cgra4ml-generated arrays in TSMC 7 nm. The 768-PE configuration closes at 1 GHz with sub-0.5 W total power (Table 6), illustrating the scalability of cgra4ml toward high-throughput ASIC targets. By varying $R{\times}C$ and memory organization from the Python frontend, users can rapidly explore power, performance, and area trade-offs across process nodes and CGRA sizes.
These results validate that the RTL emitted by cgra4ml is not tied to any single foundry or node, but can be retargeted across a wide spectrum of technologies. The ability to support both bleeding-edge 7nm and legacy 65nm nodes makes cgra4ml suitable for deployment in academic tapeouts, low-cost SoCs, and commercial research settings alike. By varying the CGRA configuration parameters from the Python frontend, users can easily explore the power, performance, and area trade-offs across process technologies. This highlights the practical utility of cgra4ml not only as a research framework but also as a prototyping and pre-silicon exploration tool for real ASIC targets. The scripts and toolflows we provide are for preliminary exploration. Due to their nature each tapeout requires an experienced physical designer to extensively tweak the scripts and flow to generate results free of DRC violations. cgra4ml allows rapid iteration on ASIC-ready CGRA accelerators, from low-power edge systems to high-performance scientific computing chips.
7. Future Work
We are expanding the framework’s capabilities to support more complex neural network architectures, including transformers, to broaden its applicability in scientific edge computing further. We are also working on migrating more and more runtime components to hardware, to parallelize their computation and minimize any CPU bottlenecks.
8. Conclusion
In this paper, we introduced cgra4ml, a modular framework for deploying large-scale neural networks in scientific edge computing. By supporting off-chip storage, diverse model architectures, and holistic ASIC/FPGA flows, cgra4ml complements existing tools like hls4ml while enabling more complex and performant designs. Its vendor-agnostic RTL generation, runtime firmware, and integrated verification infrastructure reduce development overhead and simplify hardware-software co-design. We believe cgra4ml offers a practical and extensible foundation for advancing high-performance scientific computing at the edge.